OpenStudy (itrymath):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). (2 points) y = - 1/3x - 14 y = - 1/3 x - 6 y = 3x - 6 y = 3x - 14

OpenStudy (itrymath):

i got y=3x-12 @Will.H @mathmate

OpenStudy (will.h):

"parallel to the line 3x - y = -3 " we need to have it in the form y = mx + b to figure out the slope -y = -3x - 3 y = 3x + 3 so we know that the slope is 3 now create a line that passes through (4,-2) y - (-2) = 3(x - 4) y + 2 = 3x - 12 y = 3x - 14

OpenStudy (itrymath):

OIOooOooOo

OpenStudy (itrymath):

i know hwta i did wrong!

OpenStudy (itrymath):

what

OpenStudy (itrymath):

i didnt subtract 2 from each side!

OpenStudy (will.h):

yeah.. you almost had it it's great that you are trying to do it DIY Nice job